I need help with my algebra homework! Please help!
Here is the problem:
Arnold needs to buy baseballs and baseball bats for his daughter's team. Baseballs cost $5 each, and baseball bats cost $80 each. Arnold doesn't want to spend more than $350. Write an inequality that represents this situation and graph the solution set. (I know how to graph, I just need to know how to solve the equation and what the equation is)
Thanks! :)

Answer by richwmiller(17219)

5a+80b<=350
we know that Arnold can't buy more than 4 bats since 4*80=320 and 5*80=400
4 bats and 6 balls
4*80+5*6
3 bats and 22 balls
3*80+5*22
2 bats and 38 balls
2*80+5*38
1 bat and 54 balls
16 balls = 1 bat
for every bat added remove 16 balls